OpenStudy (anonymous):

can some one help/ show me (understand) how to get to the answer? x + 2 + 3x + 5 + x + 3 + 2x

OpenStudy (anonymous):

Add all terms containing 'x' and the numbers together. By the ways, shouldn't it be something like x + 2 + 3x +5 + x + 3 + 2x = 0 ?

OpenStudy (anonymous):

no thats not have they give it to me in the book. I know what the answer is cause im suppose to simplify it. I just dont understand how to get the answer or how to do any of this.

OpenStudy (anonymous):

I just dont get Pre Algebra or regular Algebra. lol and it really makes me feel dumb

OpenStudy (anonymous):

3x + 2x + x + x + 2 + 5 + 3 Just add them together. That's the simplified form I suppose..

OpenStudy (anonymous):

It is not a quadratic equation, so I fail to see why that would be the case, Fall12. No, this is merely an expression. The key is to view all of the 'x'-containing figures as compatible terms, meaning it is permissible to add, subtract, divide, multiply, etc. them, and the non-Xes as compatible terms. Now, when it comes to this particular example, the way to do it is the following: First, look at all of the terms with an 'x' in them. These are all compatible with each other. In your case, you have x+3x+x+2x. These all add up to 7x, don't they? Precisely. Next, you look at the terms which are merely ordinary numbers without any variables, in your case: 2+5+3, which, added together, yields 10, no? Now, you combine the two. What do you end up with? 7x + 10. And that is how it's done. Math doesn't have to be hard if one tries to think logically and in steps. :)

OpenStudy (anonymous):

yes thats what i end up with

OpenStudy (anonymous):

Yeah, so we have the final expression then. Just add the like terms together.. @Veritas: I'm listening..

OpenStudy (anonymous):

I guess im just confused on how we got a 7x out of it. how did it equal to it.

OpenStudy (anonymous):

x + 2 + 3x + 5 + x + 3 + 2x x + 3x + x + 2x + 2 + 5 + 3 7x + 10

OpenStudy (anonymous):

Just add the like terms together :)

OpenStudy (anonymous):

i am really lost 2+3+5+2+3=15

OpenStudy (anonymous):

When looking solely at the variables, the ones containing X'es, we found that they add up to 7x. (x + 3x + x + 2x) + (2 + 5 + 3) = 7x + 10

OpenStudy (anonymous):

i dont get the 7 that is what i am saying

OpenStudy (anonymous):

You add the ones having 'X' in with them and the ones that don't, SEPARATELY.

OpenStudy (anonymous):

The trick is to ONLY look at the figures containing the X'es. 7x+2 = 9, for example, would be incorrect, because 7x means seven multiplied by an unknown.

OpenStudy (anonymous):

Whereas 9 simply has a numerical value of 9.

OpenStudy (anonymous):

so it would be (4)x's+the 3 that equals 7 and then do the 2+3+5=10 and thats how i get the answer

OpenStudy (anonymous):

Exactly! Think of the X-containing numbers as apples, and of the ordinary numbers as oranges. You can't add them together, because they're two different things! You instead have to add them separately. :)

OpenStudy (anonymous):

Very good.

OpenStudy (anonymous):

ok thank you both very much.

OpenStudy (anonymous):

Have fun!